A car initially traveling at 69.7 mi/h, slows to rest in 9.1 s. What is the car's acceleration?


1
Expert's answer
2021-04-13T06:42:04-0400

"0=v-at,\implies"

"a=\frac vt,"

"a=\frac{31.16}{9.1}=3.4~\frac ms."